# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ IV.3 Resolvent form and Dyson-like equation
Instead of establishing a practical scheme for density–potential inversion, this section will be devoted to an interesting reformulation of the basic relation $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ that becomes possible for a regularized functional. Starting from Eq. (18) for the universal density functional $F^{\lambda}(x)$ and using Eq. (14) we get
@@
$\nabla F_{\varepsilon}^{\lambda}=-\varepsilon^{-1}J+\varepsilon^{-1}J\circ\partial E^{\lambda}\circ(-\nabla F_{\varepsilon}).$ (77)
@@
$\partial E_{\varepsilon}^{\lambda}=\partial E^{\lambda}-\varepsilon J^{-1}$ (78)
which is just Eq. (58) again and shows another path to arrive at Eq. (77). In our setting of MY applied to DFT, the mapping $\varepsilon^{-1}J$, that gets defined by the topology of the space $X$ itself, gives the ‘bare’ connection between densities and potential. Compared to this, $\nabla F_{\varepsilon}^{\lambda}$ is a ‘dressed’ way how to connect the quantities, where the dressing occurs through the presence of particles the potential can interact with. Lastly, $\partial E^{\lambda}$ derives from the functional $E^{\lambda}(v)$, the energy of the matter system. This last element introduces a possible ambiguity into the formulation that needs to be considered if degeneracy occurs and $\partial E^{\lambda}(v)$ is a multi-valued set. Nevertheless, this reformulation allows for an interesting alternative viewpoint that links it to Green’s function methods and diagrammatic expansions.
@@ V Regularized Kohn–Sham theory
We already noted in Section II.5 that the requirement of having all involved densities simultaneously interacting and non-interacting $v$-representable puts a severe limitation on the possibility of a rigorous formulation of the Kohn–Sham method. This limitation can also influence numerical applications of the method, since it means that basis-set limits are ill-defined. The MY regularized functional $F^{\lambda}_{\varepsilon}(x)$, with the interaction strength now again noted as $\lambda$, offers a way how to avoid representability issues since the functional is differentiable and every $x\in X$ becomes representable through $\nabla F^{\lambda}_{\varepsilon}(x)$. The corresponding regularized Kohn–Sham method was first formulated in Kvaal et al. [1] for a Hilbert-space setting and later extended to Banach spaces [14]. In this section, we not only give its basic formulation, but also show how usual exchange–correlation can be translated to the new method. Finally, we repeat the main steps of a convergence proof for the method with an additional damping step. As usual, we require Assumption 1 but have to move to the Hilbert-space setting (Assumption 3) and even finite-dimensional spaces for the convergence proof. Since in the KS method we ought to stay connected to the Schrödinger equation and related KS equation, also Assumption 4 ($F^{\lambda}=F_{\mathrm{DM}}^{\lambda}$) is necessary.
@@ V.1 Iteration scheme
We will build directly on the iteration scheme introduced in Section II.5 but, with the help of MY regularization, can completely avoid the critical issues within the usual formulation of the KS method. As a reminder, we were able to introduce a well-defined exchange–correlation potential $v_{\mathrm{xc}}(\rho)$ only if $\rho$ is simultaneously interacting and non-interacting $v$-representable, i.e., at every density $\rho$ that occurs within the KS iteration both subdifferentials $\partial F^{1}(\rho)$ and $\partial F^{0}(\rho)$ must be non-empty. From what we learned about MY regularization, this can be directly guaranteed if instead we switch to the regularized functionals $F^{1}_{\varepsilon}(x)$ and $F^{0}_{\varepsilon}(x)$. However, we then need to make sense of the resulting regularized KS iteration procedure. So let us define in exact analogy to Eq. (40) for any $x\in X$ with $v=-\nabla F^{1}_{\varepsilon}(x)$ and $v_{\mathrm{s}}=-\nabla F^{0}_{\varepsilon}(x)$,
$v_{\mathrm{xc},\varepsilon}(x)$ $=v_{\mathrm{s}}-v-v_{\mathrm{H}}(x)$ (79)
$=\nabla F^{1}_{\varepsilon}(x)-\nabla F^{0}_{\varepsilon}(x)-v_{\mathrm{H}}(x)$
$=\nabla(\underbrace{F^{1}_{\varepsilon}(x)-F^{0}_{\varepsilon}(x)-E_{\rm H}(x)}_{\textstyle E_{\mathrm{xc},\varepsilon}(x)}).$
We now write $x$ instead of $\rho$ to highlight that this is an arbitrary element (see Section III.3) in $X$ that does not need to have the properties of an actual density. We can now define the same, self-consistent iteration procedure as in Eqs. (44) and (45), where we can use Eq. (58) (which is equivalent to Eq. (18)) for the superdifferential of the $E^{0}_{\varepsilon}$ functional. Let the external potential that acts on the fully interacting system be $v_{\mathrm{o}}+v$ (as before, the offset $v_{\mathrm{o}}$ will be included into the functionals), then iterate starting from an initial guess,
@@
In solving for the next $x_{i+1}$ in Eq. (81), we rely on exactly the same non-interacting Schrödinger equation that allows an orbital construction as in the usual KS procedure, we just have to shift the resulting density by $\varepsilon J^{-1}(v_{i+1})$ as the density–potential mixing dictates. If the procedure converges (see Section V.3 for a convergence proof that applies if $X$ is finite-dimensional and an additional damping step is added) we get
@@
and the resulting $x$ can then be mapped to the ground-state density with the proximal map according to Eq. (59). But now an important difference to the usual KS procedure appears that was already apparent in the formulation of the regularized KS iteration: The method aims at a shared $x\in X$ for the full and auxiliary system, and not at the internal ground-state density. This means, when $v_{\mathrm{o}}+v_{\mathrm{s}}$ is applied to the auxiliary system we get the same $x=\rho_{\mathrm{s}}-\varepsilon J^{-1}(v_{\mathrm{s}})$ as when $v_{\mathrm{o}}+v$ is applied to the full system. Therefore, in general $\rho\neq\rho_{\mathrm{s}}$ and instead we have the ground-state densities
@@
That it must hold $\rho,\rho_{\mathrm{s}}\in D$ serves as a consistency check for the method or can be used as a posteriori corrections. To get an expression for the ground-state energy, we can rely on the analogous expression to Eq. (48),
$E^{1}_{\varepsilon}(v)=E^{0}_{\varepsilon}(v)+E_{\mathrm{xc},\varepsilon}(x)-E_{\rm H}(x)-\langle v_{\mathrm{xc},\varepsilon}(x),x\rangle,$ (85)
@@ V.2 Transforming exchange–correlation potentials
While this reformulation of the KS iteration scheme with the help of MY regularization is conceptually already very appealing, the question remains how it relates back to the unregularized setting. More concretely, we want to know how $v_{\mathrm{xc},\varepsilon}(x)$ relates to $v_{\mathrm{xc}}(\rho)$, provided no $v$-representability issues arise at $\rho$. For this we take Eqs. (83) and (84) and combine them into
$v_{\mathrm{xc},\varepsilon}(x)=\varepsilon^{-1}J(\rho_{\mathrm{s}}-x)-\varepsilon^{-1}J(\rho-x)-v_{\mathrm{H}}(\rho).$ (87)
Note that if $X$ is a Hilbert space then $J$ is a linear mapping and the expression above can be simplified. Within the regularized KS iteration we already have access to $\rho_{\mathrm{s}}$ in every step as an element of $\partial E^{0}(v)$, so we just need the corresponding $\rho\in\partial E^{1}(v)$. Since we of course want to avoid solving the full problem, the $v_{\mathrm{xc}}(\rho)$ from standard KS-DFT (or any approximation to it) enters the picture, since it allows getting the same density from the auxiliary system by adding just the usual exchange–correlation potential, $\rho\in\partial E^{0}(v+v_{\mathrm{H}}(\rho)+v_{\mathrm{xc}}(\rho))$. Together with $v=\varepsilon^{-1}J(\rho-x)$, we can derive a self-consistent iteration that works similar to KS itself,
@@ V.3 Convergence proof
Now that a rigorous formulation of the KS method has been achieved, the question of convergence of the method to the desired ground-state density, when using the exact exchange–correlation functional, remains to be answered. First results on convergence of the usual KS method (Section II.5) required a $v$-representability assumption and only showed convergence in energy [69; 70]. With the help of MY regularization the problem of $v$-representability could be avoided and by the additional inclusion of a damping step a full convergence proof was achieved at least for a setting where $X$ is a (finite-dimensional) Euclidean space [71; 72; 20]. This setting means that Assumption 3 holds, and we slightly generalize the result to any finite-dimensional Hilbert space $X$ in the following presentation.
Implementations of KS regularly come with additional steps like damping that facilitate convergence. Such a damping step must also be included here, which means that in Eq. (81) not the full step from $x_{i}$ to $x_{i+1}$, but one with a reduced step length is applied. This step length will be such that one reaches exactly the apex of the regularization parabola that is tangentially aligned to $F^{1}_{\varepsilon}+v$ at $x_{i}$, as illustrated in Figure 2. Yet, it must be stressed that such a choice of step-length is not practical, since it requires knowledge about the correct functional. Its sole purpose here is to be able to establish a mathematical convergence proof. Different schemes with an a priori or adaptive choice of step length can be constructed as well [50; 14; 73]. We also note on a strong parallel to the proximal-point algorithm [74], where the minimum of $f(x)$ is found by iterating $x_{i+1}=\mathop{\mathrm{prox}}_{\varepsilon f}(x_{i})$. By the definition of the proximal map, Eq. (16), this corresponds to jumping to the apex of the full regularization function $\varepsilon\phi(x-x_{i})$ attached to $x_{i}$ in every step. The important difference to the regularized Kohn–Sham method is that the method fixes the step direction in the previous step (by solving the non-interacting system with the exchange–correlation potential) and then jumps to the apex of the regularization parabola restricted to this direction.
@@
In order to show convergence in densities and potentials, a compactness argument must be used. The revised proof [72] shows boundedness of the sequence $(x_{i})_{i}$ simply from the normalization condition in a finite-dimensional Hilbert space $X$. This means $(x_{i})_{i}$ has (possibly many) convergent subsequences of which we exemplary take $x_{\alpha(i)}\to z$. Since we are in a Hilbert-space setting (Assumption 3), property ix holds and $\nabla F^{\lambda}_{\varepsilon}$ is Lipschitz continuous. Now the first step Eq. (89) asserts
@@
so by continuity $v_{\alpha(i)+1}\to v+\nabla F^{1}_{\varepsilon}(z)-\nabla F^{0}_{\varepsilon}(z)$ and the subsequence is bounded. By the same argument as before also $x_{\alpha(i)+1}^{\prime}\in\partial E^{0}_{\varepsilon}(v_{\alpha(i)+1})$ is bounded and so we really have $\lim_{i\to\infty}(-\nabla F^{1}_{\varepsilon}(x_{\alpha(i)}))=-\nabla F^{1}_{\varepsilon}(z)=v$ from Eq. (98). This already implies $\lim_{i\to\infty}v_{\alpha(i)+1}=-\nabla F^{0}_{\varepsilon}(z)$, so we have found a potential for the non-interacting KS system reproducing $z\in\partial E^{1}_{\varepsilon}(v)$. Taking a different subsequence $x_{\beta(i)}$ can then yield another element from $\partial E^{1}_{\varepsilon}(v)$ as well as its corresponding KS potential.
@@ VI Auto-regularization in mean-field theories
In certain, specialized settings, like the Hartree-approximation to DFT or DFT with induced magnetic fields (Maxwell–Schrödinger DFT), the special form of the additional energy terms leads to automatic MY regularization. Interestingly, in the first case this is not on the side of the universal density functional, as in Eq. (49), but affects the ground-state energy $E^{\lambda}(v)$. In the case of Maxwell–Schrödinger DFT we can apply Legendre–Fenchel transformations to two different arguments and thus have the possibility to define four different functionals from convex conjugation, in some analogy to the potentials of thermodynamics that are also linked by Legendre transformations (the one-dimensional version of a Legendre–Fenchel transform). There, it turns out, that two of those functionals can even be linked to a certain form of double MY regularization called Lasry–Lions regularization.

### v2

@@ IV.3 Resolvent form and Dyson-like equation
Instead of establishing a practical scheme for density–potential inversion, this section will be devoted to an interesting reformulation of the basic relation $v=-\nabla F_{\varepsilon}^{\lambda}(x)$ that becomes possible for a regularized functional. Starting from Eq. (18) for the universal density functional $F^{\lambda}$ and using Eq. (14) we get
@@
$\nabla F_{\varepsilon}^{\lambda}=-\varepsilon^{-1}J+\varepsilon^{-1}J\circ\partial E^{\lambda}\circ(-\nabla F_{\varepsilon}^{\lambda}).$ (77)
@@
$\partial E_{\varepsilon}^{\lambda}=\partial E^{\lambda}-\varepsilon J^{-1},$ (78)
which is just Eq. (58) again and shows another path to arrive at Eq. (77). In our setting of MY applied to DFT, the mapping $\varepsilon^{-1}J$, that gets defined by the topology of the space $X$ itself, gives the ‘bare’ connection between densities and potential. Compared to this, $\nabla F_{\varepsilon}^{\lambda}$ is a ‘dressed’ way how to connect the quantities, where the dressing occurs through the presence of particles the potential can interact with. Lastly, $\partial E^{\lambda}$ derives from the functional $E^{\lambda}(v)$, the energy of the matter system. This last element introduces a possible ambiguity into the formulation that needs to be considered if degeneracy occurs and $\partial E^{\lambda}(v)$ is a multi-valued set. Nevertheless, this reformulation allows for an interesting alternative viewpoint that links it to Green’s function methods and diagrammatic expansions. We hope that this can serve as a starting point for future investigations, either for concrete approximations to the potential $\nabla F_{\varepsilon}^{\lambda}$ or into how to effectively link this back to the corresponding unregularized form $\partial F^{\lambda}$.
@@ V Regularized Kohn–Sham theory
We already noted in Section II.5 that the requirement of having all involved densities simultaneously interacting and non-interacting $v$-representable puts a severe limitation on the possibility of a rigorous formulation of the Kohn–Sham method. This limitation can also influence numerical applications of the method, since it means that basis-set limits are ill-defined. The MY regularized functional $F^{\lambda}_{\varepsilon}(x)$ offers a way how to avoid representability issues since the functional is differentiable and every $x\in X$ becomes representable through $\nabla F^{\lambda}_{\varepsilon}(x)$. The corresponding regularized Kohn–Sham method was first formulated in Kvaal et al. [54] for a Hilbert-space setting and later extended to Banach spaces [58]. In this section, we not only give its basic formulation, but also show how usual exchange–correlation can be translated to the new method. Finally, we repeat the main steps of a convergence proof for the method with an additional damping step. As usual, we require Assumption 1 but have to move to the Hilbert-space setting (Assumption 3) and even finite-dimensional spaces for the convergence proof. Since in the KS method we ought to stay connected to the Schrödinger equation and related KS equation, also Assumption 4 ($F^{\lambda}=F_{\mathrm{DM}}^{\lambda}$) is necessary.
@@ V.1 Iteration scheme
We will build directly on the iteration scheme introduced in Section II.5 but, with the help of MY regularization, can completely avoid the critical issues within the usual formulation of the KS method. As a reminder, we were able to introduce a well-defined exchange–correlation potential $v_{\mathrm{xc}}(\rho)$ only if $\rho$ is simultaneously interacting and non-interacting $v$-representable, i.e., at every density $\rho$ that occurs within the KS iteration both subdifferentials $\partial F^{1}(\rho)$ and $\partial F^{0}(\rho)$ must be non-empty. From what we learned about MY regularization, this can be directly guaranteed if instead we switch to the regularized functionals $F^{1}_{\varepsilon}(x)$ and $F^{0}_{\varepsilon}(x)$. However, we then need to make sense of the resulting regularized KS iteration procedure. So let us define in exact analogy to Eq. (40) for any $y\in X$ with $v(y)=-\nabla F^{1}_{\varepsilon}(y)$ and $v_{\mathrm{s}}(y)=-\nabla F^{0}_{\varepsilon}(y)$,
$v_{\mathrm{xc},\varepsilon}(y)$ $=v_{\mathrm{s}}(y)-v(y)-v_{\mathrm{H}}(y)$ (79)
$=\nabla F^{1}_{\varepsilon}(y)-\nabla F^{0}_{\varepsilon}(y)-v_{\mathrm{H}}(y)$
$=\nabla(\underbrace{F^{1}_{\varepsilon}(y)-F^{0}_{\varepsilon}(y)-E_{\rm H}(y)}_{\textstyle E_{\mathrm{xc},\varepsilon}(y)}).$
We now write $y$ instead of $\rho$ to highlight that this is an arbitrary element in $X$ that does not need to have the properties of an actual density (see also Section III.3). We can now define the same, self-consistent iteration procedure as in Eqs. (44) and (45), where we can use Eq. (58) (which is equivalent to Eq. (18)) for the superdifferential of the $E^{0}_{\varepsilon}$ functional. Let the external potential that acts on the fully interacting system now be fixed as $v_{\mathrm{off}}+v$ with $v=-\nabla F^{1}_{\varepsilon}(x)$ (as before, the offset $v_{\mathrm{off}}$ will be included into the functionals), then iterate starting from an initial guess,
@@
In solving for the next $x_{i+1}$ in Eq. (81), we rely on exactly the same non-interacting Schrödinger equation that allows an orbital construction as in the usual KS procedure, we just have to shift the resulting density by $\varepsilon J^{-1}(v_{i+1})$ as the density–potential mixing dictates. If the procedure converges (see Section V.3 for a convergence proof that applies to finite-dimensional $X$ if an additional damping step is added), we get
@@
and the resulting $x$ can then be mapped to the ground-state density with the proximal map according to Eq. (59). But now an important difference to the usual KS procedure appears that was already apparent in the formulation of the regularized KS iteration: The method aims at a shared $x\in X$ for the full and auxiliary system, and not at the internal ground-state density. This means that when $v_{\mathrm{off}}+v_{\mathrm{s}}$ is applied to the auxiliary system we get the same $x=\rho_{\mathrm{s}}-\varepsilon J^{-1}(v_{\mathrm{s}})$ as when $v_{\mathrm{off}}+v$ is applied to the full system. Therefore, in general $\rho\neq\rho_{\mathrm{s}}$ and instead we have the ground-state densities
@@
That it must hold $\rho,\rho_{\mathrm{s}}\in D$ serves as a consistency check for the method or can be used as an a posteriori correction. To get an expression for the ground-state energy, we can rely on the analogous expression to Eq. (48),
$E^{1}_{\varepsilon}(v)=E^{0}_{\varepsilon}(v_{\mathrm{s}})+E_{\mathrm{xc},\varepsilon}(x)-E_{\rm H}(x)-\langle v_{\mathrm{xc},\varepsilon}(x),x\rangle,$ (85)
@@ V.2 Transforming exchange–correlation potentials
While this reformulation of the KS iteration scheme with the help of MY regularization is conceptually already very appealing, the question remains how it relates back to the unregularized setting. Having again $v=-\nabla F^{1}_{\varepsilon}(x)$, in the first KS step that is Eq. (80) we rely on an exchange–correlation potential for the regularized scheme at every step $x_{i}$ for which so far only the defining relation Eq. (79) was given. More concretely, we want to know how $v_{\mathrm{xc},\varepsilon}$ relates to the usual $v_{\mathrm{xc}}$ or any approximation to it. For this we take Eqs. (83) and (84) and combine them into
$v_{\mathrm{xc},\varepsilon}(x)=\varepsilon^{-1}J(\rho_{\mathrm{s}}-x)-\varepsilon^{-1}J(\rho-x)-v_{\mathrm{H}}(x).$ (87)
Note that if $X$ is a Hilbert space then $J$ is a linear mapping and the expression above can be simplified. Within the regularized KS iteration we already have access to $\rho_{\mathrm{s}}$ in every step as an element of $\partial E^{0}(v)$, so we just need the corresponding $\rho\in\partial E^{1}(v)$ in order to evaluate the above equation. Since we of course want to avoid solving the full problem, the $v_{\mathrm{xc}}$ from standard KS-DFT (or any approximation to it) enters the picture, since it allows getting the same density from the auxiliary system by adding just the usual exchange–correlation potential, $\rho\in\partial E^{0}(v+v_{\mathrm{H}}(\rho)+v_{\mathrm{xc}}(\rho))$. Together with $v=\varepsilon^{-1}J(\rho-x)$, we can derive a self-consistent iteration that works similar to KS itself,
@@ V.3 Convergence proof
Now that a rigorous formulation of the KS method has been achieved, the question of convergence of the method to the desired ground-state density, when using the exact exchange–correlation functional, remains to be answered. First results on convergence of the usual KS method (Section II.5) required a $v$-representability assumption and only showed convergence in energy [107; 108]. With the help of MY regularization the problem of $v$-representability could be avoided and by the additional inclusion of a damping step a full convergence proof was achieved at least for a setting where $X$ is a (finite-dimensional) Euclidean space [74; 75; 55]. This setting means that Assumption 3 holds, and we slightly generalize the result to any finite-dimensional Hilbert space $X$ in the following presentation.
Implementations of the KS method regularly come with additional steps like damping that facilitate convergence. Such a damping step must also be included here, which means that in Eq. (81) not the full step from $x_{i}$ to $x_{i+1}$, but one with a reduced step length is applied. This step length will be such that one reaches exactly the apex of the regularization parabola that is tangentially aligned to $F^{1}_{\varepsilon}+v$ at $x_{i}$, as illustrated in Figure 3. Yet, it must be stressed that such a choice of step-length does not yield a practical guideline, since it requires knowledge about the fully interacting functional. Its sole purpose here is to be able to establish a mathematical convergence proof. An actual implementation of the KS method must decide about the step length from the available information from the previous steps. Different schemes with an a priori or adaptive choice of step length can be constructed as well [23; 58; 56] and their convergence rates have been compared [56; 74]. We also note on a strong parallel to the proximal-point algorithm [71], where the minimum of $f(x)$ is found by iterating $x_{i+1}=\mathop{\mathrm{prox}}_{\varepsilon f}(x_{i})$. By the definition of the proximal map, Eq. (16), this corresponds to jumping to the apex of the full regularization function $\varepsilon\phi(x-x_{i})$ attached to $x_{i}$ in every step. The important difference to the regularized Kohn–Sham method is that the method fixes the step direction in the previous step (by solving the non-interacting system with the exchange–correlation potential) and then jumps to the apex of the regularization parabola restricted to this direction.
@@
In order to show convergence in densities and potentials, a compactness argument must be used. The revised proof [75] uses the fact that $F^{1}_{\varepsilon}(x)$ increases like $\|x\|^{2}_{X}$ asymptotically outside of its effective domain $D$ as a result of MY regularization. For $X$ finite-dimensional, which critically enters here, all norms are equivalent and thus $D$ is bounded since it only includes normalized densities. Since $e_{i}=F_{\varepsilon}^{1}(x_{i})+\langle v,x_{i}\rangle<e_{1}$ from the result that the $e_{i}$ are strictly decreasing, we arrive at $(x_{i})_{i}$ bounded. Again, the finite dimensionality of $X$ is critical in the next step, because it guarantees that a bounded sequence always has convergent subsequences. This means $(x_{i})_{i}$ has (possibly infinitely many) convergent subsequences of which we exemplary take $x_{\alpha(i)}\to z$. Since we are in a Hilbert-space setting (Assumption 3), property ix holds and $\nabla F^{\lambda}_{\varepsilon}$ is Lipschitz continuous. Now the first step Eq. (89) asserts
@@
so by continuity $v_{\alpha(i)+1}\to v+\nabla F^{1}_{\varepsilon}(z)-\nabla F^{0}_{\varepsilon}(z)$ and the subsequence is bounded. By the same argument as before also the sequence given by $x_{\alpha(i)+1}^{\prime}\in\partial E^{0}_{\varepsilon}(v_{\alpha(i)+1})$ is bounded which means $\|x_{\alpha(i)+1}^{\prime}-x_{\alpha}(i)\|_{X}$ is bounded and so we really have $\lim_{i\to\infty}(-\nabla F^{1}_{\varepsilon}(x_{\alpha(i)}))=-\nabla F^{1}_{\varepsilon}(z)=v$ from Eq. (98). This already implies $\lim_{i\to\infty}v_{\alpha(i)+1}=-\nabla F^{0}_{\varepsilon}(z)$, so we have found a potential for the non-interacting KS system reproducing $z\in\partial E^{1}_{\varepsilon}(v)$. Taking a different subsequence $x_{\beta(i)}$ can then yield another element from $\partial E^{1}_{\varepsilon}(v)$ as well as its corresponding KS potential.
@@ VI Auto-regularization in mean-field theories
In certain, specialized settings, like the Hartree-approximation to DFT or DFT with induced magnetic fields (Maxwell–Schrödinger DFT), the special form of the additional energy terms leads to automatic MY regularization. Interestingly, in the first case this is not on the side of the universal density functional, as in Eq. (49), but affects the ground-state energy $E^{\lambda}(v)$. In the case of Maxwell–Schrödinger DFT we can apply Legendre–Fenchel transformations to two different arguments and thus have the possibility to define four different functionals from convex conjugation, in some analogy to the potentials of thermodynamics that are also linked by Legendre–Fenchel transformations. There, it turns out that two of those functionals can even be linked to a certain form of double MY regularization called Lasry–Lions regularization.